The liquid used in automobile cooling systems is prepared by dissolving ethylene glycol (HOCH2CH2OH) in water. Ethylene glycol has a molar mass of 62.07 g/mol and a density of 1.115 g/mL at 50.0°C.

Calculate the vapor pressure at 50°C of a coolant solution that is 54.0:46.0 ethylene glycol-to-water by volume. At 50.0°C, the density of water is 0.9880 g/mL, and its vapor pressure is 92 torr. The vapor pressure of ethylene glycol is less than 1 torr at 50.0°C.
Answer #1

100 ml of solution contains 54 ml of ethylene glycol and 46 ml of water.

Mass of ethylene glycol = 54 ml × 1.115 g/ ml = 60.21 g

Moles of ethylene glycol = mass/ molar mass = 60.21 g/ (62.07 g/ mole) = 0.97 mole

Mass of water = 46 ml × 0.9880 g/ ml = 45.448 g

Moles of water = mass / molar mass of water = 45.448 g / (18 g/ mole) = 2.525mole

Mole fraction of water in solution = moles of water / (moles of water + moles of ethylene glycol) = 2.525/(0.97 + 2.525) = 0.722

Vapor pressure of coolant solution = 92 torr × 0.722 = 66.47 torr